A cyclist leaves his training base for a morning workout, riding at the rate of 16 mph. One and one-half hours later, his support staff leaves the base in a car going 40 mph in the same direction. How long will it take the support staff to catch up with the cyclist?
Answer by checkley79(3341)

D=RT
D=16T
D=40*1.5
16T=40*1.5
16T=60
T=60/16
T=3.75 HOURS THEY WILL MEET.